# Conditional Text-to-Image Generation with Reference Guidance

### v1

@@ 1 Introduction
Recent developments in text-to-image models have yielded groundbreaking achievements, showcasing an unprecedented capability in translating natural language descriptions into visually compelling depictions. Despite the remarkable capabilities of generating images given general textual instructions, off-the-shelf text-to-image models usually struggle with precisely generating specific subjects at a high standard. For example, Stable Diffusion (SD) [45] is known for English misspelling.
@@
Several research works have sought to address this issue by substituting conventional CLIP text encoders with a more powerful language model such as T5 [43]. As demonstrated by DeepFloyd [2], Imagen [49] and eDiff-I [9], large language models enhance both sample fidelity and image-text alignment. While scaling up language models can enhance language comprehension, it comes at a substantial cost of computation and model training. Another line of research strives to resolve this challenge with additional conditions that exhibit spatial correlation with the input. TextDiffuser [12] utilized a segmentation mask predicted by a layout generation module as an additional condition. In a similar vein, GlyphDraw [38] proposed using text renderings of target keywords to condition the model for generating Chinese scene-text images. However, these methods demonstrate limited capability in generating images that are beyond the vocabulary of the text encoder they are trained with. For instance, modern text-to-image models universally employ English language models [42] for prompt embedding, hindering their generalization to image generation with text from a different language (Conditional Text-to-Image Generation with Reference Guidance, middle) or accurate rendition of non-text subjects such as logos (Conditional Text-to-Image Generation with Reference Guidance, right).
@@
To resolve the aforementioned shortcoming of text-to-image diffusion models, we propose RefDiffuser , an approach that uses reference images as an additional source of condition, improving the generation results of particular subjects of interest with the guidance of explicit visual reference of their appearance. Our method is built upon the SD [45], which adopts an UNet for progressive denoising in a learned latent space of images. Instead of training an independent network branch or new layers from scratch to process the new visual condition, we directly encode the reference image into the same latent space using the VAE. More specifically, we augment the first layer of SD UNet to simultaneously accept the noisy image latent and the reference latent as the inputs. We then finetune the UNet to learn the natural blending of the visual references and the textual instruction to generate high-fidelity images following both conditions. Considering the large scales of the state-of-the-art diffusion models, tuning full models for diverse tasks can be prohibitive in terms of both parameter size and memory footprints. We resort to the low-rank adaptation of the SD models and develop a series of small-scale expert plug-ins, each containing only up to 28.55M parameters. Training objectives and auxiliary networks are customized for various applications, including English scene-text generation, multi-lingual scene-text generation, and logo-image generation. We present comprehensive discussions on the sampling strategies and showcase the distinct impacts of both conditions in different denoising time steps, ahieving an English scene-text, multi-lingual scene-text, and logo generation accuracy of 61.73%, 46.88%, 44.07 with high prompt fidelity, outperforming the baselines.
@@ 2 Related Works
Diffusion models. Diffusion models Surpassing the prior family of generative models such as GAN [17, 7, 41, 26] VAE [44, 31, 54, 58], diffusion models [53, 21, 56] have demonstrated remarkable capabilities in generating images with both high quality and diversity, either in pixel space [21, 55, 27] or a learned latent space [45, 10, 59]. With the help of advanced pretrained language models [42, 43] and sampling techniques [55, 20], text-to-image diffusion models [45, 49, 39, 11, 40, 9, 61] show unprecedented results on image generation following textual instruction. They generate high-resolution images by either operating in latent space [45, 18] or using cascaded models [49, 9] to progressively scale up resolution. Motivated by the known issue of misspelling for almost all the public text-to-image diffusion models, several works have been proposed to improve text drawing by additional conditions such as masks [12] and glyph [38]. However, the model design customized for a particular language [12] prevents them from generalizing to general visual references. Recent research has been striving to improve the generation quality of particular subjects by model tuning. Specifically, model customization [48, 30, 60, 13, 52, 32, 8] focuses on a transfer learning approach that tunes model parameters to fix new concepts given examples. Textual inversion [16] learns a new concept by learning a new word vector. These methods fit well to particular subjects but usually fail to generalize to other similar ones. A line of research works [65, 33, 63, 34, 15] such as ControlNets [65] have been proposed to enhance the controllability of the generation using a new input condition.
@@
Visual text generation. Despite the progress in generating images in high-quality images, the existing generative model has been noted to generate visually inaccurate images (i.e., misspellings). Several research studies have been proposed to mitigate this issue [36, 49, 2, 38, 12]. One line of research has shown that the visual accuracy of the text renderings can be improved by deploying a language model of larger capacity [2, 49, 38, 62]. Imagen [49] has demonstrated that encoding text prompts with the generic text encoder T5 [43] pretrained on text-only corpora can improve both the fidelity and image-text alignment [49]. To evaluate the accuracy of the generated visual text images, several papers have constructed their own benchmark sets [36, 38, 12] using off-the-shelf OCR models [23, 35, 28, 64, 24].
Figure 2: Overview of the model training. The model receives a reference image as a condition in addition to the text prompt. The reference image and binary mask jointly guide the location, and the reference image further guides the shape and uncurated appearance of the generation target. An auxiliary network receives the pooled features from the target regions and is trained online with each expert plugin to encourage accurate drawing of the target subjects.
@@ 3 Method
The proposed RefDiffuser generates images conditioned on both a text prompt and a reference image. The image condition gives the model uncurated visual reference for the generation targets, such as character shapes in scene-text image generation. We train the model to blend the reference concept naturally into the generated images without violating the text conditions. This reference condition empowers the model to produce contents that the original diffusion model fails to generate precisely, and can even help extend the generation to concepts not included in the language model vocabulary. Our proposed model is constructed on the foundation of the pre-trained Stable Diffusion model [45]. We employ the low-rank adaptation technique [22] to fine-tune the model for specific tasks, such as scene-text image generation, and develop a series of small-scale plug-ins, each seamlessly converting a Stable Diffusion model into an expert model for different applications with references. In the subsequent section, we elaborate on the model architecture, training methodologies, and inference procedures in detail.
@@ 3.1 Overview
The proposed RefDiffuser builds upon the foundation of the pre-trained SD models [45] with a pre-trained CLIP [42] text encoder to encode text prompts, and is trained to incorporate additional image conditions. Training additional network branches to process the reference images introduces considerable cost [65]. Exploiting the fact that the references reside in the same image modality, we directly encode the reference images into the same latent space using the VAE [29] encoder of SD, and augment the input layer of the diffusion UNet [47], expanding the inputs with the additional reference latent and a spatial mask indicating the location of the targets of interest in the reference image.
@@
where $\mathbf{s}\in\mathbb{R}^{c\times h\times w}$ is a VAE encoded representation of the spatial reference image, and $t\sim[1,T]$ is the step within the $T$ total time steps. The location mask $\mathbf{m}\in\mathbb{R}^{h\times w}$ in (Equation 1) is a binary array indicating the location and desired shape of the reference objects, providing extra guidance on how to deform the uncurated target appearance specified by the reference image.The binary location mask $\mathbf{m}$ is downsampled to the same spatial size as the latent. The mask $\mathbf{m}$ and reference latent $\mathbf{s}$ remain constant across all time steps and are not injected with any noise in contrast to $\mathbf{z}$. With this input composition, we now have $\mathbf{x}^{t}\in\mathbb{R}^{(2c+1)\times h\times w}$. If not otherwise specified, we follow the standard training of SD-v1.5, which has $c=4$ and $h=w=64$ when generating images at $512\times 512$. The reconstruction loss of diffusion training is defined as:
$\mathcal{L}_{\text{diff}}=\sum_{i=1}^{N}||\epsilon_{i}-\epsilon_{\theta}(\mathbf{x}^{t}_{i},\mathbf{c}_{i},t)||^{2},$ (2)
@@
Directly tuning $\theta$ introduces a high cost in terms of both parameter sizes and memory footprints, hindering the broad applications of models on diverse tasks. Therefore, we resort to the low-rank adaptation method [22] and develop expert plug-ins for different applications, each containing only $\sim 28.55$M parameters. We present details regarding the low-rank adaptation in Section A.1.
@@ 3.2 Example Plugin 1: Text Image Generation
In scene-text image generation, we train our model to generate high-quality images following the text prompt that describes a scene with English words involved in the scene, e.g., “A dog holding a sign saying ‘Hello World’.”. SD models are known to struggle with correct English spelling, so we provide the model with an additional reference of the characters to draw in the image and use rendered text as our reference image for the text image generation as visualized in Figure 2. The standard OCR detectors [23, 35] can be used to locate text within the images, and we use OCR detection results to render the reference for our model. Specifically, given a natural scene image and the OCR labels, we generate the reference image by rendering the text to the corresponding region on a blank canvas. We also generate the binary location mask where positive values indicate the precise size and shape of the desired text to draw in the image. With the reference image encoded using the SD VAE for $\mathbf{s}$ and binary mask $\mathbf{m}$ resized to the same spatial size as the latent, we construct the model input as in (Equation 1).
@@
$\mathcal{L}_{\text{recog}}=\frac{-1}{NL}\sum_{i=1}^{N}\sum_{j=1}^{L}y_{i,j}\log\boldsymbol{\psi_{\text{recog}}}(\hat{\mathbf{z}}^{0}_{i},\mathbf{l}^{i}_{j}),$ (3)
where $N$ represents the batch size, enumerated by $i$, and $L$ denotes the maximum length of a word set, enumerated by $j$. We use $y_{i,j}\in\mathbb{R}^{|C|}$ as the one-hot labels and the location information of the $j$-th word in the $i$-th image, respectively. The set of characters is denoted as $C$. Note that this network can be discarded after training, and the final model consists of only the original SD UNet and a lightweight expert plug-in.
The final learning objective of the diffusion network, $\mathcal{L}_{\epsilon}$, combines the reconstruction loss and the recognition loss through a weighted sum as:
$\mathcal{L_{\epsilon}}=\lambda_{\text{recog}}\mathcal{L_{\text{recog}}}+\lambda_{\text{diff}}\mathcal{L_{\text{diff}}}.$ (4)
In contrast to [38], our method does not strictly require the target text word to appear in the text prompt. Consequently, the model can draw the text either based on both the text prompt and the visual reference or rely on the visual reference only, offering additional flexibility.
Multi-Lingual Text Image Generation. The flexibility of the conditions allows us to extend the generation to concepts beyond the vocabulary of the language model SD trained with. To show this, we develop an expert plug-in for multi-lingual text (MLT) image generation that covers multiple languages. Due to the difficulties in collecting large-scale datasets for MLT generation, we first pretrain the diffusion and recognition network on the large-scale English OCR dataset [12], then proceed with the fine-tuning. The overall learning process is mostly the same as the English text image generation training, except that the size of the alphabet is expanded to cover all the target languages simultaneously. We resort to public OCR datasets [14] for high-quality samples. Due to insufficient images in the existing MLT OCR datasets, we augment training data with synthesized images by manually collecting text-free background images [5], fonts in different scripts, and text corpora of each language, then rendering text words in random regions inside the background images (e.g., Figure 3, Greek, Russian, Thai.). We then use a merged dataset with both real images from MLT OCR datasets and the synthesized image. Directly training the model with the mixed dataset compromises the quality of the generated images, as they can exhibit noticeable artifacts inherited from the synthetic images. To remedy this, we introduce an additional scaler $\alpha$ to scale the diffusion reconstruction loss:
@@
$+\sum_{i\notin\text{synth}}^{N}||\epsilon_{i}-\epsilon_{\theta}(\mathbf{x}^{t}_{i},\mathbf{c}_{i},t)||^{2}.$
We scale down $\alpha$ for the synthetic images to prevent the diffusion model from overfitting to the synthetic artifacts when learning to compose the scene and exploit the synthetic data mainly for improving the spelling correctness with the help of the recognition network. We observe this simple loss scaling significantly improves the generated image quality. Loss turns related to the recognition network are identical to the ones used in the English scene-text generation.
@@ 3.3 Example Plugin 2: Logo Image Generation
We show that the proposed framework can be generalized to non-text objects and use logo image generation as an example to show the versatility of the proposed conditional image generation with reference guidance. For the reference images of logo image generation, we paste the standard logo we collected from the Internet to a blank canvas and create the reference images. We visualize one example reference image in Figure 2. Similar to the MLT case, we synthesize logo images to augment the training set (Figure 4), and we deploy an auxiliary network to classify logos. Specifically, we take the predicted denoised latent features at timestep 00, $\hat{\mathbf{z}}^{0}$, and apply the RoIAlign [19] to pool the region. Then, similar to the text image generation, we replace it with the uncorrupted latent $\mathbf{z}^{0}$ for the same reason. Once the RoI features are obtained, we feed them to the auxiliary network $\psi_{\text{logo}}(\cdot)$ to do the classification:
$\mathcal{L}_{\text{logo}}=\frac{-1}{K}\sum_{k=1}^{K}y_{k}\log\psi_{\text{logo}}(\mathbf{z}^{0}_{k}),$ (6)
where $K$ denotes the total number of RoI instances in the batch enumerated by $k$, $y_{k}\in\mathbb{R}^{|M|}$ is the one-hot labels. The overall learning objective can be formulated as follows:
@@
Although we train the model with a closed set of logos, the model can generalize to the ones that are unseen during training. We show visual examples to validate this in Figure 7, and Figure F in Appendix.
@@ 3.4 Scheduled Classifier-Free Guidance
Classifier-free guidance [20] has become the standard technique for sampling with conditional diffusion models. It substantially improves the generated image quality and alignment with the condition prompt by extrapolating the predicted $\epsilon$ toward the direction of the condition:
$\vspace{-3mm}\hat{\epsilon_{\theta}}^{t}:=\epsilon_{\theta}(\mathbf{z}^{t},\mathbf{c},t)+\omega(\epsilon_{\theta}(\mathbf{z}^{t},\mathbf{c},t)-\epsilon_{\theta}(\mathbf{z}^{t},\varnothing,t)),\vspace{-2mm}$ (8)
@@
Augmented Classifier-Free Guidance. With the additional reference condition included, we opt to reformulate the CFG as below to enable more detailed guidance of each conditions,
$\begin{split}\hat{\epsilon}^{t}&=\underbrace{\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)}_{\text{Unconditional}}\\ &\underbrace{+\omega_{\text{prompt}}\Big{[}\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\mathbf{c},t)-\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)\Big{]}}_{\text{Caption Guidance}}\\ &+\underbrace{\omega_{\text{ref}}\Big{[}\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\textbf{s},\textbf{m}),\varnothing,t)-\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)\Big{]}}_{\text{Reference Guidance}}\\ &+\underbrace{\omega_{\text{all}}\Big{[}\epsilon(\text{concat}(\mathbf{z}^{t},\textbf{s},\textbf{m}),\mathbf{c},t)-\epsilon(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)\Big{]}}_{\text{All Guidance}},\end{split}$ (9)
@@ 4 Experiments
4.1 Training Datasets
We explain the dataset that we use for training. Additional details of the training sets are available in Section A.5.
Training Dataset for English. For training English plug-in models, we first train the model with MARIO-10M [12] and further fine-tune the model by including samples from LAION-Aesthetic [51] and additional aesthetic text images (AText) that we collect (Figure 5).
Training Dataset for Multi-Lingual Text. For the MLT plug-in, we finetune the English plug-in model trained with MARIO-10M using the merged set of real images from ICDAR2019 [14], and synthetic images that we generated.
Training Dataset for Logos. Similar to the MLT plug-in, we also finetune the English plug-in model trained with MARIO-10M using the merged set of real images from the merged set of FlickrLogos-32 [25] and Logos in the wild (LITW) [57], and the synthetic images that we generated.
4.2 Benchmark
Benchmark for English Generation. We adopt the MARIO-Eval [12] benchmark to assess the model performance in generating English scene-text images. The MARIO-Eval benchmark contains 5,414 prompts. This benchmark is composed of six different subsets, each a subset of multiple benchmarks, including DrawBenchText [36], DrawTextCreative [36], ChineseDrawText [38], and Mario-10M test set [12]. Following the protocol [12], we exclude images generated from ChineseDrawText and DrawBenchText, hence 5,000 images are used for measuring FID.
Benchmark for Multi-Lingual Generation. Due to the lack of pre-existing benchmarks for MLT image generation, we developed a new benchmark, for model evaluation on nine different languages. We create 25 prompt templates, such as “A raccoon holding a paper saying words.”. We exclude the non-English target words from the prompts, as they cannot be processed by the text encoder. Each prompt is designed to include up to three words in the image. Except for English, we allocate the rest of eight languages a set of 200 prompts, and 400 for English, yielding 2,500 prompts in total. For computing FID, we construct a set of images by randomly sampling a subset of the ICDAR2019 dataset.
Benchmark for Logo Generation. We additionally construct a benchmark for logo image generation. Specifically, we create 25 template prompts, e.g., “A building displaying a [KEYWORD] sign.”, where [KEYWORD] here is the name of the logo. We obtain 150 logos in total, which is the union of the class of FlickrLogos-32 dataset [25] and LITW dataset, yielding 1,500 prompts in total. For computing FID, we construct an image set sampled from the merged set of FlickrLogos-32 [46] and LITW [57] dataset.
4.3 Evaluation Metrics
We use Fréchet inception distance (FID) to assess the image quality, and we evaluate the semantic alignment by computing the CLIP Score between the text prompts and the generated images and measure the aesthetic scores of the generated images obtained from a pretrained model [51].
Text images. To measure the accuracy of the generated text, we use Accuracy and F1-score. We apply different OCR Engines for English generation tasks and MLT image generation tasks. We use Microsoft Read API for English, and Google Cloud Vision API for MLT image generation tasks. We consider a detected text correct if it is exactly matched with a keyword and the F-1 score is defined as the harmonic mean of precision and recall.
Logo images. To measure the accuracy of the logo image generation, we train a Mask-RCNN [19] on the merged set of FlickrLogos-32 [46] and LITW dataset [57]. We report the Accuracy and F1-score of the logo detection model applied to an image set generated by models.
4.4 Implementation Details.
We first fine-tune the pre-trained Stable Diffusion (SD) V2.1 [45] on the MARIO-10M dataset [12]. For English image generation, we further train the model using the AText dataset we collect. For MLT and Logo image generation, we train the model with merged sets of synthetic images and real images. We set batch size=10 for each of the 8 GPUs. We use AdamW [37] with a learning rate of 1e-4 to tune the parameter. For MLT image generation, we use a character set containing 847 characters from the nine languages. We use a rank of 32 for all LoRA tuning if not otherwise specified. We use 100 steps of DDIM [55] sampling. For all the experiments, we set the recognition loss weight $\lambda$ as 0.025.
@@ 4.3 Evaluation Metrics
${}^{\text{\textendash}}$ 22.50 0.3037
Table 1: Evaluation results of Russian generation. ${}^{\text{\textendash}}$ denotes the model trained without seeing any Russian images during training. The accuracy is denoted in [%].

### v2

@@ Abstract
Recent developments in text-to-image models have yielded groundbreaking achievements, showcasing an unprecedented capability in translating natural language descriptions into visually compelling depictions. Despite the remarkable capabilities of generating images given general textual instructions, off-the-shelf text-to-image models usually struggle with precisely generating specific subjects at a high standard. For example, Stable Diffusion (SD) [SD] is known for English misspelling.
@@
Several research works have sought to address this issue by substituting conventional CLIP text encoders with a more powerful language model, such as T5 [t5]. As demonstrated by DeepFloyd [deepfloyd], Imagen [imagen] and eDiff-I [ediffi], large language models enhance both sample fidelity and image-text alignment. While scaling up language models can enhance language comprehension, it comes at a substantial cost of computation and model training. Another line of research strives to resolve this challenge with additional conditions that exhibit spatial correlation with the input. TextDiffuser [textdiffuser] utilized a segmentation mask predicted by a layout generation module as an additional condition. In a similar vein, GlyphDraw [glyphdraw] proposed using text renderings of target keywords to condition the model for generating Chinese scene-text images. However, these methods demonstrate limited capability in generating images that are beyond the vocabulary of the text encoder they are trained with. For instance, modern text-to-image models universally employ English language models [clip] for prompt embedding, hindering their generalization to image generation with text from a different language (Figure 1, middle) or accurate rendition of non-text subjects such as logos (Figure 1, right).
@@
To resolve the aforementioned shortcoming of text-to-image diffusion models, we propose RefDiffuser , an approach that uses reference images as an additional source of condition, improving the generation results of particular subjects of interest with the guidance of explicit visual reference of their appearance. Our method is built upon the SD [SD], which adopts an UNet for progressive denoising in a learned latent space of images. Instead of training an independent network branch or new layers from scratch to process the new visual condition, we directly encode the reference image into the same latent space using the VAE. More specifically, we augment the first layer of SD UNet to simultaneously accept the noisy image latent and the reference latent as the inputs. We then finetune the UNet to learn the natural blending of the visual references and the textual instruction to generate high-fidelity images following both conditions. Considering the large scales of the state-of-the-art diffusion models, tuning full models for diverse tasks can be prohibitive in terms of both parameter size and memory footprints. We resort to the low-rank adaptation of the SD models and develop a series of small-scale expert plug-ins, each containing only up to 28.55M parameters. Training objectives and auxiliary networks are customized for various applications, including English scene-text generation, multi-lingual scene-text generation, and logo-image generation. We present comprehensive discussions on the sampling strategies and showcase the distinct impacts of both conditions in different denoising time steps, achieving an English scene-text, multi-lingual scene-text, and logo generation accuracy of 61.73%, 46.88%, 44.07 with high prompt fidelity, outperforming the baselines.
@@
Diffusion models. Diffusion models Surpassing the prior family of generative models such as GAN [gan, wgan, dcgan, stylegan] VAE [vae1, vae2, cvae, vqvae], diffusion models [sohl2015deep, ho_diffusion, song2020score] have demonstrated remarkable capabilities in generating images with both high quality and diversity, either in pixel space [ho_diffusion, ddim, edm] or a learned latent space [SD, unleashing, bld]. With the help of advanced pretrained language models [clip, t5] and sampling techniques [ddim, cfg], text-to-image diffusion models [SD, imagen, glide, chang2023muse, podell2023sdxl, ediffi, raphael] show unprecedented results on image generation following textual instruction. They generate high-resolution images by either operating in latent space [SD, gu2022vector] or using cascaded models [imagen, ediffi] to progressively scale up resolution. Motivated by the known issue of misspelling for almost all the public text-to-image diffusion models, several works have been proposed to improve text drawing by additional conditions such as masks [textdiffuser] and glyph [glyphdraw]. However, the model design customized for a particular language [textdiffuser] prevents them from generalizing to general visual references. Recent research has been striving to improve the generation quality of particular subjects by model tuning. Specifically, model customization [dreambooth, multi, elite, suti, instantbooth, blip, break_a_scene] focuses on a transfer learning approach that tunes model parameters to fix new concepts given examples. Textual inversion [ti] learns a new concept by learning a new word vector. These methods fit well to particular subjects but usually fail to generalize to other similar ones. A line of research works [controlnet, controlnet2, reco, gligen, make_a_scene] such as ControlNets [controlnet] have been proposed to enhance the controllability of the generation using a new input condition.
@@
Visual text generation. Despite the progress in generating images in high-quality images, the existing generative model has been noted to generate visually inaccurate images (i.e., misspellings). Several research studies have been proposed to mitigate this issue [drawtext, imagen, deepfloyd, glyphdraw, textdiffuser]. One line of research has shown that the visual accuracy of the text renderings can be improved by deploying a language model of larger capacity [deepfloyd, imagen, glyphdraw, glyphcontrol]. Imagen [imagen] has demonstrated that encoding text prompts with the generic text encoder T5 [t5] pretrained on text-only corpora can improve both the fidelity and image-text alignment [imagen]. To evaluate the accuracy of the generated visual text images, several papers have constructed their own benchmark sets [drawtext, glyphdraw, textdiffuser] using off-the-shelf OCR models [swinsptter, maskspotter, units, deepsolo, synspotter].
While some existing works rely on text renderings to improve English and multi-lingual generation [textdiffuser, glyphcontrol, anytext], or focus on personalized object generation [dreambooth, blip], our work is distinguished by introducing a single, unified framework that bridges these capabilities. By leveraging lightweight, task-specific expert plugins, our method achieves high fidelity in producing accurate multi-lingual text as well as complex brand logos, thereby validating its broad applicability.
Figure 2: Overview of the model training. Expert plugin modules are trained on each dataset. The input and reference conditions are encoded using a VAE, and the mask is resized and concatenated with the reference. An additional convolutional layer (Input Conv2) processes these concatenated features, and its output is combined with the existing input convolution layers. The regional features are obtained from denoised latents, and an auxiliary network receives them. The auxiliary network is trained online with each expert plugin to encourage accurate drawing of the target subjects. The Input Conv2 and the LoRA parameters are updated during training.
@@
The proposed RefDiffuser generates images conditioned on both a text prompt and a reference image. The image condition gives the model an uncurated visual reference for the generation targets, such as character shapes in scene-text image generation. We train the model to blend the reference concept naturally into the generated images without violating the text conditions. This reference condition empowers the model to produce contents that the original diffusion model fails to generate precisely, and can even help extend the generation to concepts not included in the language model vocabulary. Our proposed model is constructed on the foundation of the pre-trained Stable Diffusion model [SD]. We employ the low-rank adaptation technique [lora] to fine-tune the model for specific tasks, such as scene-text image generation, and develop a series of small-scale plug-ins, each seamlessly converting a Stable Diffusion model into an expert model for different applications with references.
@@
The proposed RefDiffuser builds upon the foundation of the pre-trained SD models [SD] with a pre-trained CLIP [clip] text encoder to encode text prompts, and is trained to incorporate additional image conditions. Training additional network branches to process the reference images introduces considerable cost [controlnet]. Exploiting the fact that the references reside in the same image modality, we directly encode the reference images into the same latent space using the VAE [vae] encoder of SD, and augment the input layer of the diffusion UNet [unet], expanding the inputs with the additional reference latent and a spatial mask indicating the location of the targets of interest in the reference image.
@@
where $\mathbf{s}\in\mathbb{R}^{c\times h\times w}$ is a VAE encoded representation of the spatial reference image, and $t\sim[1,T]$ is the step within the $T$ total time steps. The location mask $\mathbf{m}\in\mathbb{R}^{h\times w}$ in (Equation 1) is a binary array indicating the location and desired shape of the reference objects. The binary location mask $\mathbf{m}$ is downsampled to the same spatial size as the latent. The mask $\mathbf{m}$ and reference latent $\mathbf{s}$ remain constant across all time steps and are not injected with any noise in contrast to $\mathbf{z}$. With this input composition, we now have $\mathbf{x}^{t}\in\mathbb{R}^{(2c+1)\times h\times w}$.
To handle the additional $c+1$ channels, we introduce an extra input convolutional layer (Input Conv2, Figure 2). Its output is added feature-wise to that of the original input convolution, and the combined feature map is then passed to the subsequent layers of the diffusion model for reconstruction. The reconstruction loss of diffusion training is defined as:
[…]
@@
Directly tuning $\theta$ introduces a high cost in terms of both parameter sizes and memory footprints. Therefore, we resort to the low-rank adaptation method [lora] and develop expert plug-ins for different applications, each containing only a small number of parameters.
@@
In scene-text image generation, our goal is to produce high-quality images that not only align with the text prompt describing a scene but also accurately generate the target text. To achieve precise text generation, we provide the model with an additional reference image in which the desired characters are pre-rendered at the corresponding location in the scene, as illustrated in Figure 2. We use OCR detection results [textdiffuser] to locate texts within the images, which can easily be obtained with the standard OCR detectors [swinsptter]. Specifically, given the OCR labels, we generate the reference image by rendering the text to the corresponding region on a blank canvas. We also generate the binary location mask, where positive values indicate the precise size and shape of the desired text to draw in the image. With the reference image encoded using the SD VAE for $\mathbf{s}$ and binary mask $\mathbf{m}$ resized to the same spatial size as the latent, we construct the model input as in (Equation 1).
To ensure the spelling accuracy of the character sequence within the mask region, we facilitate the training with an online-trained lightweight text recognition network. We detail the network architecture of the recognition network in Section A.1. Given the noise estimation predicted by the diffusion network, we reconstruct the denoised latent at timestep 0, denoted as $\hat{\mathbf{z}}^{0}$, from the corrupted latent $\mathbf{z}^{t}$ [ho_diffusion]. We employ RoIAlign [mask_rcnn] to extract text regions into a fixed size, and these pooled regions are fed into an online-trained text recognition network $\boldsymbol{\psi_{\text{recog}}}(\cdot)$ to compute the text recognition loss. The recognition loss is computed as,
$\small\textbf{r}_{k}$ $=\text{RoIAlign}(\hat{\mathbf{z}}^{0},\mathbf{B}_{k})$ (3)
$\textbf{o}_{k}$ $=\psi_{\text{recog}}(\textbf{r}_{k})$ (4)
$\mathcal{L}_{\text{recog}}$ $=\frac{-1}{L}\sum_{j=1}^{L}y_{j}\log\textbf{o}_{k,j},$ (5)
where $\textbf{B}_{k}$ is the bounding box label for $k$-th region in the image, $\textbf{o}_{k}$ is the network output. $L$ denotes the length of the $k$-th word and $j$ is an enumerator for each character in the word. Here, $\textbf{o}_{k,j}\in\mathbb{R}^{|C|}$ is the predicted probability over the character set at each position $j$, and $y_{j}\in\mathbb{R}^{|C|}$ is the corresponding one-hot label. We provide more details of the loss in Section A.1. Note that, unlike TextDiffuser, our method only requires word-level information, instead of character-level segmentation supervision [textdiffuser]. This network can be discarded after training, and the final model consists of only the original SD UNet and a lightweight expert plug-in. The final learning objective of the diffusion network, $\mathcal{L}_{\epsilon}$, combines the reconstruction loss and the recognition loss through a weighted sum as:
$\small\mathcal{L_{\epsilon}}=\lambda_{\text{recog}}\mathcal{L_{\text{recog}}}+\lambda_{\text{diff}}\mathcal{L_{\text{diff}}}.$ (6)
Multi-Lingual Text Image Generation. The flexibility of the conditions allows us to extend the generation to concepts beyond the vocabulary of the language model SD trained with. To show this, we develop an expert plug-in for multi-lingual text (MLT) image generation that covers multiple languages. Due to the difficulties in collecting large-scale datasets for MLT generation, we first pretrain the plugin module and recognition network on the large-scale English OCR dataset [textdiffuser], then proceed with the fine-tuning. The overall learning process is mostly the same as the English text image generation training, except that the size of the alphabet is expanded to cover all the target languages simultaneously. Due to insufficient images in the existing MLT OCR datasets, we augment training data with synthesized images by manually collecting text-free background images [pexels], fonts in different scripts, and text corpora of each language, then rendering text words in random regions inside the background images (e.g., Figure 3, Greek, Russian, Thai.). We then use a merged dataset with both real images from MLT OCR datasets [icdar2019] and the synthesized image. Directly training the model with the mixed dataset compromises the quality of the generated images, as they can exhibit noticeable artifacts inherited from the synthetic images. To remedy this, we introduce an additional scaler $\alpha$ to scale the diffusion reconstruction loss:
@@
We scale down $\alpha$ for the synthetic images to prevent the diffusion model from overfitting to the synthetic artifacts when learning to compose the scene and exploit the synthetic data mainly for improving the spelling correctness with the help of the recognition network. We observe that this simple loss scaling effectively improves the generated image quality. Loss terms related to the recognition network are identical to the ones used in the English scene-text generation.
@@
We further show that the proposed framework can be generalized to non-text objects and use logo image generation as an example to show the versatility of the proposed conditional image generation with reference guidance.
For the reference images of logo image generation, we paste the standard logo we collected from the Internet onto a blank canvas and create the reference images. We visualize one example reference image in Figure 2. Similar to the MLT case, we synthesize logo images to augment the training set (Figure 4), and we deploy an auxiliary network to classify logos. To compute the loss, we predict the denoised latent features at timestep 0, $\hat{\mathbf{z}}^{0}$, and apply the RoIAlign [mask_rcnn] to extract the region $\textbf{r}_{k}=\text{RoIAlign}(\hat{\textbf{z}}^{0},\textbf{B}_{k})$. The RoI features are then provided to the auxiliary network $\psi_{\text{logo}}(\cdot)$ to do the classification:
$\mathcal{L}_{\text{logo}}=\frac{-1}{K}\sum_{k=1}^{K}y_{k}\log\psi_{\text{logo}}(\textbf{r}_{k}),$ (8)
where $K$ denotes the total number of RoI instances in the batch enumerated by $k$, $y_{k}\in\mathbb{R}^{|M|}$ is the one-hot labels for $|M|$ logos we use for training. The overall learning objective can be formulated as follows:
@@
Although we train the model with a closed set of $|M|$ logos, the model can generalize to the ones that are unseen during training. We show visual examples to validate this in Figure 7, and Figure I in the Appendix.
@@
Classifier-free guidance [cfg] has become the standard technique for sampling with conditional diffusion models. It substantially improves the generated image quality and alignment with the condition prompt by extrapolating the predicted $\epsilon$ toward the direction of the condition:
$\small\hat{\epsilon_{\theta}}^{t}:=\epsilon_{\theta}(\mathbf{z}^{t},\mathbf{c},t)+\omega(\epsilon_{\theta}(\mathbf{z}^{t},\mathbf{c},t)-\epsilon_{\theta}(\mathbf{z}^{t},\varnothing,t)),$ (10)
@@
Augmented Classifier-Free Guidance. With the additional reference condition included, we opt to reformulate the CFG as below to enable more detailed guidance of each condition,
$\small\begin{split}\hat{\epsilon}^{t}&=\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)\\ &+\omega_{\text{prompt}}\Big[\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\mathbf{c},t)-\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)\Big]\\ &+\omega_{\text{ref}}\Big[\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\textbf{s},\textbf{m}),\varnothing,t)-\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)\Big]\\ &+\omega_{\text{all}}\Big[\epsilon(\text{concat}(\mathbf{z}^{t},\textbf{s},\textbf{m}),\mathbf{c},t)-\epsilon(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)\Big],\end{split}$ (11)
@@
[…]